If Microsoft decides to invest 3 million dollars, and believes it will have sales of 100k in the first year and after that sales will increase by 25% per year for the next 3 years. Also, after those 3 years sales will then grow at a constant 3% rate forever. Market risk premium of 4.4% Should I make this investment? Why or why not?
Answer #1

Let's say, risk free rate=1.54% (1 year T bill rate)

Required rate of return=Risk free rate+Market Risk Premium=1.54%+4.4%=5.94%

First year sales=$100k and 4th year sales=100*(1+25%)^3=$195.31K

Terminal Cash flow or 5th year cash flow= 4th year cash flow*(1+growth rate)/(Required raet of return-growth rate)

=195.31*(1+3%)/(5.94%-3%)=6842.493

Now, below is the calculation of NPV:

1 Year Present Value of cash flow Cash flow(in $1000) (Cash flow/1.0594^year) (3,000.00) (3,000.00) 100.00 94.39 125.00 111.3

The project has NPV of $2619.84 which is greater than zero.

Hence, you should invest into the project.